Minimum area of circle which touches the parabola’s y = x2 + 1 and y2 = x – 1 is



Grade:10

1 Answers

Amit Saxena
10 years ago

B)

 

            parabola''s y = x2 + 1 and x = y2 + 1 are symmetrical about y = x

 

            \         tangent of point A is parallel to y = x